# On the gcd graphs over polynomial rings and related topics

### v1

@@ 4. Bipartite property
A graph $G=(V,E)$ is called a bipartite graph if $V$ can be partitioned into two disjoint sets $V_{1},V_{2}$ such that for every edge $(u,v)\in E$ either $u\in V_{1},v\in V_{2}$ or $u\in V_{2},v\in V_{1}$. Bipartite graphs model various real-life situations such as job assignments, resource allocation, stable marriage matching, gene-disease association, and much more. Additionally, from a theoretical point of view, bipartite graphs often provide a good test for theorems and algorithms on graphs. We refer the reader to [2] for some further discussions on this topic.
The goal of this section is to classify all $G_{f}(D)$ which is bipartite. We remark that while our focus is on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$, everything we discuss in this section applies to gcd-graphs over $\operatorname{\mathbb{Z}}$ as well (see 4.5). To the best of our knowledge, even over $\operatorname{\mathbb{Z}}$, this topic has not been explored in the literature. Therefore, in addition to studying this problem for its own merits, we also hope to address a gap in the literature.
@@
    $I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 0\pmod{2}\right\},$
    $I_{1}=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 1\pmod{2}\right\},$
@@
  1. (1)
    Let us consider the first part of this lemma. Because $G_{f}(D)$ is bipartite, we can write
@@
    where $A,B$ are two independent sets in $G_{f}(D).$ Without loss of generality, we can assume that $0\in A$. We claim that $A=I$. By the proof of [4, Proposition 2.6], we know that $A$ is a subgroup of $(R,+)$ of index $2$ By definition, if $h_{i}\in S_{D}$, then $(0,h_{i})\in E(G_{f}(D))$. Because $A$ and $B$ are disjoint independent set in $G_{f}(D)$, $h_{i}\in B$. Furthermore, since $A$ is a subgroup of index $2$ in $R$, $h_{i}+h_{j}\in A$ for all $h_{i},h_{j}\in S_{D}.$ Consequently, we must have $I\subseteq A.$
@@
  2. (2)
    Suppose that $G_{f}(\{1\})$ is connected. We will show that $I$ is an ideal in $R$ as well. The idea is similar to the proof of 3.2 so we will be brief. Specifically, we note that for each $s\in R^{\times}$, $sI=I.$ Since $\langle R^{\times}\rangle=R$, this shows that $aI\subseteq I$ for all $a\in R.$ We conclude that $I$ is an ideal in $R$.
  3. (3)
    Finally, part $(3)$ follows from Equation Eq. 3.1.
@@
If $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$, we know that $G_{f}(\{1\})$ is connected by 3.3. Furthermore, $\operatorname{\mathbb{F}}_{q}[x]/f$ has no ideal of index $2$ By 4.1, we conclude that $G_{f}(D)$ is not bipartite. ∎
@@
Finally, let us consider the trickest case where $x(x+1)\mid f.$
@@
$B=\{g\in R|g(0)\neq g(1)\}=x+A.$
We can check that if $a_{1},a_{2}\in A$ then $\gcd(a_{1}-a_{2},x(x+1))\in\{1,x(x+1)\}.$ By definition, $(a_{1},a_{2})\not\in E(G_{f}(D)).$ This shows that $A$ is an independent set in $G_{f}(D)$. Similarly, $B$ is an independent set in $G_{f}(D)$ as well. We conclude that $G_{f}(D)$ is not bipartite.
@@
$I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}n_{i}\equiv 0\pmod{2}\right\},$
is an independent set in $G_{f}(D).$ Furthermore, by part $(3)$ of 3.7, we know that $x(x+1)R\subset I.$ Because 00 is not connected to any node in $I$, we must have $x(x+1)R\cap S_{D}=\emptyset$. Consequently $x(x+1)\not\in\overline{D}.$ Suppose to the contrary that $|\overline{D}|\geq 3$. We then have $\overline{D}=\{1,x,x+1\}.$ This implies that $G_{x(x+1)}(\overline{D})$ is the complete graph $K_{4}.$ In $K_{4}$, there is a path of length $2$ from 00 to any other vertices. From this property, we conclude that for each $g\in R$, we can write
@@
for some $h_{1},h_{2}\in S_{D}$. We then conclude that $g-h_{1}-h_{2}\in x(x+1)R\subset I$. Since $h_{1}+h_{2}\in I$, we conclude that $g\in I$ as well. This shows that $I=R$, which is a contradiction. ∎
@@
By an almost identical argument, we can show that the gcd graph $G_{n}(D)$ over $\operatorname{\mathbb{Z}}$ with $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and $\gcd(d_{1},d_{2},\ldots,d_{k})=1$, is bipartite if and only if $2\mid n$ and $2\nmid d_{i}$ for all $1\leq i\leq k.$
5. Primeness of gcd graphs
@@
The concept of a homogeneous set appears in various branches of mathematics (see [6, 8, 12, 25] for some concrete examples). One of our main motivations comes from the fact that homogeneous sets allow us to decompose a network into a multilevel network of smaller graphs. From both theoretical and computational perspectives, such a decomposition is crucial for understanding the dynamics of multilevel networks (see [5, 13, 17, 25]). In [8], in collaboration with some other graph theorists, we completely classify prime unitary Cayley graphs on finite commutative rings. In this section, we study the problem for the gcd graph $G_{f}(D).$ We remark that for a graph $G$, a connected component of $G$ (or of its complement) is a homogeneous set in $G$. Therefore, while studying the primeness of $G$, it is safe to assume that $G$ is connected and anti-connected. For this reason, we will assume throughout this section that $G_{f}(D)$ is both connected and anti-connected (we refer the reader to Section 3 for precise conditions for these properties to hold).
@@
Let $a\in(\operatorname{\mathbb{F}}_{q}[t]/f)^{\times}.$ Let $m_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/f$ be the multiplication by $a$ map. Then $m_{a}$ induces an automorphism on $G_{f}(D).$
@@
Since $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$, $m_{a}$ is an automorphism of $(\operatorname{\mathbb{F}}_{q}[x]/f,+)$. Furthermore, $m_{a}$ preserves $S_{D}$; i.e $aS_{D}=S_{D}.$ As a result, $m_{a}$ is an automorphism of $G_{f}(D).$
@@
    There exists a non-trivial ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $I$ is a homogeneous set in $G_{f}(D).$
@@
Clearly $(2)$ implies $(1)$. Let us now prove $(1)$ implies $(2)$. The proof that we discuss here is quite similar to the one that we gave for [8, Theorem 4.1] and 3.2. Because $G_{f}(D)$ is not prime, we can find a maximal homogenous set $H$ containing $0$ By [8, Theorem 3.4], we know that $H$ is a subgroup of $\operatorname{\mathbb{F}}_{q}[x]/f$. We claim that it is an ideal as well. By 5.1, we conclude that $aH=H$ for all $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}.$ Furthermore, since $G_{f}(\{1\})$ is connected, $aH\subseteq H$ for all $a\in\operatorname{\mathbb{F}}_{q}[x]/f$ as well. This shows that $H$ is an ideal in $\operatorname{\mathbb{F}}_{q}[x]/f.$
@@
We remark that the relation $a-bt_{1}=a_{1}c$ and the fact that $\gcd(a,b)=1$ imply that $\gcd(a_{1},b)=1.$ By the Chinese remainder theorem, we can find $t_{2}$ such that $\gcd(a_{1}-bt_{2},\frac{f}{c})=1.$
@@
Let $g$ be a divisor of $f$. Let $I_{g}$ be the ideal in $R=\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ The induced graph on $I_{g}$ is isomorphic to $G_{f/g}(D_{g})$ where
$D_{g}=\left\{\frac{f_{i}}{g}|f_{i}\in D,g\mid f_{i}\right\}.$
@@
Let $g\mid f$ be a divisor of $f$ and $I$ the ideal in $\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ Let
$D_{1}=\left\{f_{i}\in D\mid g\nmid f_{i}\right\},\quad D_{2}=\left\{f_{i}\in D|g\mid f_{i}\right\}.$
@@
$\overline{D_{1}}=\left\{\gcd(f_{i},g)\mid f_{i}\in D_{1}\right\},\quad\widetilde{D_{2}}=\left\{\frac{f_{i}}{g}\mid f_{i}\in D_{2}\right\}.$
@@
    $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ where $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ is the canonical projection map.
Furthermore, if one of the above equivalent conditions holds, $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(D_{1})*G_{f/g}(\widetilde{D_{2}}).$
@@
First, let us show that $(2)$ implies $(1)$. In other words, suppose that $g$ satisfies the condition that $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$. We claim that the ideal $I$ generated by $g$ is a homogeneous set in $G_{f}(D).$ In fact, let $a\not\in I$ be an arbitrary element in $\operatorname{\mathbb{F}}_{q}[x]/f$ and suppose that $a$ is adjacent to an element in $I$. By a translation, we can assume that $a$ is adjacent to 00 in $G_{f}(D).$ We claim that $a$ is adjacent to all elements in $I$ as well. Let $gt$ be an element in $I$. We need to show that
@@
Suppose this is not the case. We can find a non-constant irreducible polynomial $h$ of such that $\gcd(h,\frac{g}{f_{i}})=1$ and $h|\frac{f}{f_{i}}.$ We then see that $\gcd(hf_{i},g)=f_{i}=\gcd(f_{i},g).$ This implies that $hf_{i}\in D$. Since $f_{i}\mid hf_{i}$ and $f_{i}\neq hf_{i}$, this leads to a contradiction.
@@
The proof of 5.6 relies crucially on the divisibility relationship between $f_{i}$ and $f_{j}.$ It seems important to study this relationship systematically. We are studying this problem in a work in progress (see [22]).
@@
While we primarily focus on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$ in this section, most statements have a straightforward analog for gcd-graphs over $\operatorname{\mathbb{Z}}.$
6. Spectrum of gcd graphs
The spectrum of gcd-graphs over $\operatorname{\mathbb{Z}}$ is described by the theory of Ramanujan sums (see [18, Section 4]), which in turn is a special case of Gauss sums (see [21]). As explained in [7], these sums are precisely values of Fekete polynomials at certain $n$-roots of unity. One might wonder whether such a similar statement for the spectrum of gcd-graphs holds in the context of function fields. While we are not able to find an analog of Fekete polynomials over function fields, the theory of Ramanujan sums does have an interesting analogy as we will explain in this section. This, however, is sufficient for us to describe explicitly the spectrum of gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x].$
@@ 6.1. Symmetric algebras
A key point in the theory of $\operatorname{\mathbb{Z}}/n$-circulant graph is the fact that the character group of $\operatorname{\mathbb{Z}}/n$ is isomorphic to $\operatorname{\mathbb{Z}}/n$:
$\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
This isomorphism can be obtained as follows. Fix a primitive $n$-root of unity $\zeta_{n}$ in $\operatorname{\mathbb{C}}$. Let $\chi_{1}\colon\operatorname{\mathbb{Z}}/n\to\operatorname{\mathbb{C}}^{\times}$ be the character defined by $\chi_{1}(m)=\zeta_{n}^{m}$ for all $m\in\operatorname{\mathbb{Z}}/n.$ For each $a\in\operatorname{\mathbb{Z}}/n$, let $\chi_{a}=\chi_{1}^{a}$ be the character of $\operatorname{\mathbb{Z}}/n$ defined by $\chi_{a}(b)=\zeta_{n}^{ab}=\chi_{1}^{a}(b).$ The following proposition is standard.
@@
The map $a\mapsto\chi_{1}^{a}$ gives an isomorphism between $\operatorname{\mathbb{Z}}/n$ and $\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
In summary, once we fix a primitive $n$-root of unity, the isomorphism $\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times})$ is obtained via the multiplicative structure on $\operatorname{\mathbb{Z}}/n$. We will use a similar approach in the function field case. We first recall the following definition.
@@
(See [20, Page 66-67]) Let $A$ be a finite dimensional commutative $\operatorname{\mathbb{F}}_{q}$-algebra. $A$ is said to be a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra if there exists an $\operatorname{\mathbb{F}}_{q}$-linear functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{q}$ such that the kernel of $\lambda$ contains no nonzero ideal of $A$. We call $\lambda$ a non-degenerate linear functional on $A.$
@@
Fix a primitive $p$-root of unity $\zeta_{p}\in\operatorname{\mathbb{C}}^{\times}$. Then, $\operatorname{\mathbb{F}}_{p}$ is (non)-cannonically a subgroup of $\operatorname{\mathbb{C}}^{\times}.$ If $A$ is an $\operatorname{\mathbb{F}}_{q}$-algebra, then $(A,+)$ is a direct sum of several copies of $\operatorname{\mathbb{F}}_{p}$. Therefore
@@
Let $A$ be a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra together with a non-degenerate functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{p}.$ For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{C}}^{\times}$ defined by $\lambda_{a}(b)=\zeta_{p}^{\lambda(ab)}.$ Then $\lambda\in\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times}).$ Furthermore, the map $a\mapsto\lambda_{a}$ gives an isomorphism between $A$ and $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})$.
We will now focus on the case $A=\operatorname{\mathbb{F}}_{q}[x]/f$. We will show that it is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra (and hence a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra as explained in 6.4). We will show this by constructing an explicit $\operatorname{\mathbb{F}}_{q}$-linear functional on $A.$ We learned about this construction in [19]. Every element $g$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ can be written uniquely in the form
$g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}.$
We define $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ by
@@
Suppose $g\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\psi(hg)=0$ for all $h\in\operatorname{\mathbb{F}}_{q}[x]/f.$ Then $g=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$
@@
Let us write $g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}$. We will prove by induction that $a_{n-k}(g)=0$ for $1\leq k\leq n.$ In fact, since $\psi(g)=0$, we know that $a_{n-1}(g)=0.$ Consequently, the statement is true for $k=1.$ Let us assume that it has been for all $1\leq k\leq m<n$. We claim that it is also true for $m+1$, namely $a_{n-m-1}(g)=0$ as well. In fact, we have
$\begin{split}x^{m-1}g&=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}+a_{n-m}(g)x^{n}+\cdots+a_{n-1}(g)x^{m+n-2}\\ &=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}.\end{split}$
Consequently
$a_{n-m-1}(g)=\psi(x^{m-1}g)=0.$

### v2

@@ 4. Bipartite property
A graph $G=(V,E)$ is called a bipartite graph if $V$ can be partitioned into two disjoint sets $V_{1},V_{2}$ such that for every edge $(u,v)\in E$ either $u\in V_{1},v\in V_{2}$ or $u\in V_{2},v\in V_{1}$. In this case, we write $V=V_{1}\sqcup V_{2}$ and we call $V_{1},V_{2}$ independent sets in $G.$
Bipartite graphs model various real-life situations such as job assignments, resource allocation, stable marriage matching, gene-disease association, and much more. Additionally, from a theoretical point of view, bipartite graphs often provide a good test for theorems and algorithms on graphs. We refer the reader to [3] for some further discussions on this topic.
[…]
@@
    $I=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 0\pmod{2}\right\}$,
    $I_{1}=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 1\pmod{2}\right\}$.
@@
(1) Let us consider the first part of this lemma. Because $G_{f}(D)$ is bipartite, we can write
@@
where $A,B$ are two independent sets in $G_{f}(D).$ Without loss of generality, we can assume that $0\in A$. We claim that $A=I$. By the proof of [5, Proposition 2.6], we know that $A$ is a subgroup of $(R,+)$ of index $2.$ By definition, if $h_{i}\in S_{D}$, then $(0,h_{i})\in E(G_{f}(D))$. Because $A$ and $B$ are disjoint independent set in $G_{f}(D)$, $h_{i}\in B$. Furthermore, since $A$ is a subgroup of index $2$ in $R$, $h_{i}+h_{j}\in A$ for all $h_{i},h_{j}\in S_{D}.$ Consequently, we must have $I\subseteq A.$
@@
(2) Suppose that $G_{f}(\{1\})$ is connected. We will show that $I$ is an ideal in $R$ as well. The idea is similar to the proof of 3.2 so we will be brief. Specifically, we note that for each $s\in R^{\times}$, $sI=I.$ Since $\langle R^{\times}\rangle=R$, this shows that $aI\subseteq I$ for all $a\in R.$ We conclude that $I$ is an ideal in $R$.
(3) Finally, part $(3)$ follows from Equation Eq. 3.1. ∎
@@
If $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$, we know that $G_{f}(\{1\})$ is connected by $3.$ Furthermore, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ has no ideal of index $2.$ By 4.1, we conclude that $G_{f}(D)$ is not bipartite. ∎
@@
Finally, let us consider the trickiest case where $x(x+1)\mid f.$
@@
$B=\{g\in R:g(0)\neq g(1)\}=x+A.$
We can check that if $a_{1},a_{2}\in A$ then $\gcd(a_{1}-a_{2},x(x+1))\in\{1,x(x+1)\}.$ By definition, $(a_{1},a_{2})\not\in E(G_{f}(D)).$ This shows that $A$ is an independent set in $G_{f}(D)$. Similarly, $B$ is an independent set in $G_{f}(D)$ as well. We conclude that $G_{f}(D)$ is bipartite.
@@
$I=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}n_{i}\equiv 0\pmod{2}\right\}$,
is an independent set in $G_{f}(D).$ Furthermore, by part $(3)$ of 3.7, we know that $x(x+1)R\subset I.$ Because 0 is not connected to any node in $I$, we must have $x(x+1)R\cap S_{D}=\emptyset$}. Consequently $x(x+1)\not\in\overline{D}.$ Suppose to the contrary that $|\overline{D}|\geq 3$. We then have $\overline{D}=\{1,x,x+1\}.$ This implies that $G_{x(x+1)}(\overline{D})$ is the complete graph $K_{4}.$ In $K_{4}$, there is a path of length $2$ from 0 to any other vertices. From this property, we conclude that for each $g\in R$, we can write
@@
for some $h_{1},h_{2}\in S_{D}$. We then conclude that $g-h_{1}-h_{2}\in x(x+1)R\subset I$. By the definition of $I$, $h_{1}+h_{2}\in I$, we conclude that $g\in I$ as well. This shows that $I=R$, which is a contradiction. ∎
@@
[…]
5. Primeness of gcd-graphs
@@
The concept of a homogeneous set appears in various branches of mathematics (see [7, 9, 13, 26] for some concrete examples). One of our main motivations comes from the fact that homogeneous sets allow us to decompose a network into a multilevel network of smaller graphs. From both theoretical and computational perspectives, such a decomposition is crucial for understanding the dynamics of multilevel networks (see [6, 14, 18, 26]). In [9], in collaboration with some other graph theorists, we completely classify prime unitary Cayley graphs on finite commutative rings. In this section, we study the problem for the gcd-graph $G_{f}(D).$ We remark that for a graph $G$, a connected component of $G$ (or of its complement) is a homogeneous set in $G$. Therefore, while studying the primeness of $G$, it is safe to assume that $G$ is connected and anti-connected. For this reason, we will assume throughout this section that $G_{f}(D)$ is both connected and anti-connected (we refer the reader to Section 3 for precise conditions for these properties to hold).
@@
Let $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}.$ Let $m_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(f)$ be map induced by the multiplication by $a$. Then $m_{a}$ induces an automorphism on $G_{f}(D).$
@@
Since $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}$, $m_{a}$ is an automorphism of $(\operatorname{\mathbb{F}}_{q}[x]/(f),+)$. Furthermore, $m_{a}$ preserves $S_{D}$; i.e $aS_{D}=S_{D}.$ As a result, $m_{a}$ is an automorphism of $G_{f}(D).$
@@
    There exists a non-trivial ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $I$ is a homogeneous set in $G_{f}(D).$
@@
Clearly $(2)$ implies $(1)$. Let us now prove $(1)$ implies $(2)$. The proof that we discuss here is quite similar to the one that we gave for [9, Theorem 4.1] and 3.2. Because $G_{f}(D)$ is not prime, we can find a maximal homogenous set $H$ containing $0.$ By [9, Theorem 3.4], we know that $H$ is a subgroup of $\operatorname{\mathbb{F}}_{q}[x]/(f)$. We claim that it is an ideal as well. By 5.1, we conclude that $aH=H$ for all $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}.$ Furthermore, since $G_{f}(\{1\})$ is connected, $aH\subseteq H$ for all $a\in\operatorname{\mathbb{F}}_{q}[x]/(f)$ as well. This shows that $H$ is an ideal in $\operatorname{\mathbb{F}}_{q}[x]/(f).$
@@
We remark that the relation $a-bt_{1}=a_{1}c$ and the fact that $\gcd(a,b)=1$ imply that $\gcd(a_{1},b)=1.$ By the Chinese remainder theorem, we can find $t_{2}$ such that $\gcd(a_{1}-bt_{2},\frac{f}{c})=1$}. ∎
@@
Let $g$ be a divisor of $f$. Let $I_{g}$ be the ideal in $R=\operatorname{\mathbb{F}}_{q}[x]/(f)$ generated by $g.$ The induced graph on $I_{g}$ is isomorphic to $G_{f/g}(D_{g})$ where
$D_{g}=\left\{\frac{f_{i}}{g}:f_{i}\in D,g\mid f_{i}\right\}.$
@@
Let $g\mid f$ be a divisor of $f$ and $I$ the ideal in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ generated by $g.$ Let
$D_{1}=\left\{f_{i}\in D:g\nmid f_{i}\right\},\quad D_{2}=\left\{f_{i}\in D:g\mid f_{i}\right\}.$
@@
$\overline{D_{1}}=\left\{\gcd(f_{i},g):f_{i}\in D_{1}\right\},\quad\widetilde{D_{2}}=\left\{\frac{f_{i}}{g}:f_{i}\in D_{2}\right\}.$
@@
    $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ where $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(g)$ is the canonical projection map.
[…]
@@
First, let us show that $(2)$ implies $(1)$. In other words, suppose that $g$ satisfies the condition that $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$. We claim that the ideal $I$ generated by $g$ is a homogeneous set in $G_{f}(D).$ In fact, let $a\not\in I$ be an arbitrary element in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ and suppose that $a$ is adjacent to an element in $I$. By a translation, we can assume that $a$ is adjacent to 0 in $G_{f}(D).$ We claim that $a$ is adjacent to all elements in $I$ as well. Let $gt$ be an element in $I$. We need to show that
@@
Suppose this is not the case. We can find a non-constant irreducible polynomial $h$ of such that $\gcd(h,\frac{g}{f_{i}})=1$ and $h\mid\frac{f}{f_{i}}.$ We then see that $\gcd(hf_{i},g)=f_{i}=\gcd(f_{i},g).$ This implies that $hf_{i}\in D$. Since $f_{i}\mid hf_{i}$ and $f_{i}\neq hf_{i}$, this leads to a contradiction.
@@
The proof of 5.6 relies crucially on the divisibility relationship between $f_{i}$ and $f_{j}.$ It seems important to study this relationship systematically. Since the completion of this paper, we have made some partial progress on this problem (see [23]).
@@
[…]
6. Spectrum of gcd-graphs
[…]
@@ 6.1. Symmetric algebras
A key point in the theory of $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph is the fact that the character group of $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is isomorphic to $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$:
$\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times}).$
[…]
@@
The map $a\mapsto\chi_{1}^{a}$ gives an isomorphism between $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ and $\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times}).$
In summary, once we fix a primitive $n$-root of unity, the isomorphism $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times})$} is obtained via the multiplicative structure on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$. We will use a similar approach in the function field case. We first recall the following definition.
@@
(See [21, Page 66-67]) Let $A$ be a finite dimensional commutative $\operatorname{\mathbb{F}}_{q}$-algebra. $A$ is said to be a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra if there exists an $\operatorname{\mathbb{F}}_{q}$-linear functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{q}$ such that the kernel of $\lambda$ contains no nonzero ideal of $A$. We call $\lambda$ a non-degenerate linear functional on $A.$
@@
[…]
@@
Let $A$ be a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra together with a non-degenerate functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{p}.$ For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{C}}^{\times}$ defined by $\lambda_{a}(b)=\zeta_{p}^{\lambda(ab)}.$ Then $\lambda_{a}\in\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times}).$ Furthermore, the map $a\mapsto\lambda_{a}$ gives an isomorphism between $A$ and $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})$.
We will now focus on the case $A=\operatorname{\mathbb{F}}_{q}[x]/(f)$. We will show that it is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra (and hence a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra as explained from 6.4). We will show this by constructing an explicit $\operatorname{\mathbb{F}}_{q}$-linear functional on $A.$ We learned about this construction from [20]. Assume $\deg f=n$. Every element $g$ in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ can be written uniquely in the form
$g=a_{0}(g)+a_{1}(g)x+\cdots+a_{n-1}(g)x^{n-1}.$
We define $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ by
@@
Suppose $g\in\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $\psi(hg)=0$ for all $h\in\operatorname{\mathbb{F}}_{q}[x]/(f).$ Then $g=0$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$
@@
Let us write $g=a_{0}(g)+a_{1}(g)x+\cdots+a_{n-1}(g)x^{n-1}$. We will prove by induction that $a_{n-k}(g)=0$ for $1\leq k\leq n.$ In fact, since $\psi(g)=0$, we know that $a_{n-1}(g)=0.$ Consequently, the statement is true for $k=1.$ Let us assume that it has been shown for all $1\leq k\leq m<n$. We claim that it is also true for $m+1$, namely $a_{n-m-1}(g)=0$ as well. In fact, we have
$\begin{split}x^{m}g&=x^{m}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}+a_{n-m}(g)x^{n}+\cdots+a_{n-1}(g)x^{m+n-1}\\ &=x^{m}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}.\end{split}$
Here, the last equation follows from the fact that $a_{n-m}=a_{n-m+1}=\ldots=a_{n-1}=0$ by the induction hypothesis. Consequently
$a_{n-m-1}(g)=\psi(x^{m}g)=0.$